OpenStudy (anonymous):

|x - 1| = |x2 -2x + 1|

OpenStudy (anonymous):

|x-1|=|x-1|^2 therefore the real solution is x=1

OpenStudy (anonymous):

You need to use the definition of absolute value and notice a couple of things. First,

OpenStudy (anonymous):

\[|u|:=\sqrt{u^2}\]and so\[\sqrt{(x-1)^2}=\sqrt{(x^2-2x+1)^2}\]Square both sides and note you can factor the right-hand side as such:\[(x-1)^2=((x-1)^2)^2\]Hence,\[(x-1)^2=(x-1)^4 \rightarrow (x-1)^2(1-(x-1)^2)=0\]Therefore, either\[(x-1)^2=0 \rightarrow x=1\]or\[1-(x-1)^2=0 \rightarrow (x-1)^2=1 \rightarrow x-1=\pm 1 \rightarrow x=1 \pm 1\]That is,\[x=0,2\]

OpenStudy (anonymous):

So there are three solutions: x=0 x=1 x=2

OpenStudy (anonymous):

This is a plot of what you're trying to do. You have two functions,\[y=|x-1|\]and\[y=|x^2-2x+1|\]and you're trying to find the points of intersection.

OpenStudy (anonymous):

THANK YOU SO SO MUCH LOKISAN AND IAMPAUL!

OpenStudy (anonymous):

np :)

OpenStudy (anonymous):

could you also help me with this , thanxs! |2x - 1| = x2

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

I need a second to do something...

OpenStudy (anonymous):

no problemo but i guess lokisan explains it miles better ;)

OpenStudy (anonymous):

:)